OpenStudy (anonymous):

write the canonical equation of the straight line that goes through the points A(0,3) and B(4, -4).

OpenStudy (anonymous):

find the slope \[M = \frac{ -4-3 }{ 4-0 }\] which equals -7/4

OpenStudy (anonymous):

you can use the points given to find the equation of the line... using the equation (y-y1)=m(x-x1) if you use point A(0,3) your equation would look like y-3=-7/4(x-0) then use algebra to put it in whatever form you want, I like point slope form because its easy to manipulate.

OpenStudy (anonymous):

would you mind telling me how will it look like in the slope form?

OpenStudy (anonymous):

y=-7/4x +3